Please help in solving the following, I have no idea where to start.
The area of a square is 81 cm^2. If the same amount is added to one dimenstion and removed from the other, the resulting rectangle has an area of 9cm^2 less than the area of the original square. How much is added and subtracted.
Thank you so much.

Answer by edjones(8007)

Let x be the amt. added and subtracted.
(9+x)(9-x)=72
81-x^2=72
-x^2=-9
x^2=9
x=3
.
Check:
(9+3)(9-3)=72 true
.
Ed